# AICO: Feature Significance Tests for Supervised Learning

### v3

@@ Abstract
The opacity of many supervised learning algorithms remains a key challenge, hindering scientific discovery and limiting broader deployment—particularly in high-stakes domains. This paper develops model- and distribution-agnostic significance tests to assess the influence of input features in any regression or classification algorithm. Our method evaluates a feature’s incremental contribution to model performance by masking its values across samples. Under the null hypothesis, the distribution of performance differences across a test set has a non-positive median. We construct a uniformly most powerful, randomized sign test for this median, yielding exact $p$-values for assessing feature significance and confidence intervals with exact coverage for estimating population-level feature importance. The approach requires minimal assumptions, avoids model retraining or auxiliary models, and remains computationally efficient even for large-scale, high-dimensional settings. Experiments on synthetic tasks validate its statistical and computational advantages, and applications to real-world data demonstrate its ability to inform high-stakes decisions with legal, economic, and regulatory implications.
@@ 1 Introduction
Supervised learning algorithms have become foundational tools across scientific and industrial domains, enabling breakthroughs and transforming key sectors. However, their opacity poses a major challenge, hindering scientific discovery and limiting adoption in high-stakes domains like finance, public policy and health care, where transparency is crucial.
This paper proposes novel testing methods for addressing the opacity of learning algorithms, focusing on the role of an algorithms’s input variables (features). Specifically, we develop a family of hypothesis tests to assess the significance of features in any supervised learning model, whether for regression or classification. The tests are model- and distribution-agnostic: they only require the ability to query the trained model—no assumptions about the model class, data distribution or model target (e.g. conditional mean, quantile) are needed. The computational efficiency of our tests is unprecedented. The procedure does not require retraining or surrogate models, making it well-suited for analyzing complex, high-dimensional algorithms that are costly to train.
The tests are based on a new approach for evaluating the impact of a feature. We quantify a feature’s incremental contribution to a user-defined score function, which measures a model’s predictive or decision-making performance on a test sample. Specifically, we compute the feature effect as the change in score when the model is given the actual value of the feature versus when the feature is masked with a baseline value. This masking approach, which we call Add-In Covariates (AICO), avoids the variance inflation (excess randomness) introduced by permutation methods and the high cost of retraining-based and surrogate-model-based strategies. The score function and masking baseline can be user-specified to align with specific interpretability goals and downstream model use cases; we discuss a number of choices and the associated tradeoffs.
@@
The feature effect measures the impact of a feature’s variability on model score; it represents the importance of a feature at the level of an individual sample. A null-feature’s effect has a distribution across a collection of test samples whose median is non-positive. We construct a uniformly most powerful, randomized sign test for this median. Randomization at the boundary of the critical region ensures the test’s exactness, i.e., its size (type I error probability) matching the user-specified nominal level (e.g. $5\%$). The test delivers exact, randomized feature $p$-values that have, conditionally on the data, a uniform distribution on an interval which can be reported in practice. It is simple, robust and permits an efficient implementation. It requires minimal assumptions, ensuring its broad validity in practice. These properties allow users to make valid inferences regarding a model’s features without relying on large-sample approximations or parametric assumptions.
@@
The median feature effect being tested is a natural measure of feature importance at the population level. We develop randomized, exact-coverage confidence intervals for the median that, together with the $p$-values, enable principled inference regarding feature influence. The availability of confidence intervals distinguishes the median from widely-used algorithmic measures of feature importance (see Section 1.1 below). Moreover, depending on the selection of the masking baseline, the median can have several useful properties. These include robustness to correlation between features, which is common in practice. Furthermore, for linear regression, the median is proportional to the square of the feature’s regression coefficient, which is the canonical measure of importance in the linear context. Finally, we provide conditions guaranteeing that a feature’s median being 0 is equivalent to it being conditionally independent of the response given the remaining features.
@@
Extensive experiments on synthetic “known truth” tasks demonstrate the statistical and computational performance of our approach in nonlinear regression and classification settings. We compare our tests with the recently developed, model-agnostic and non-asymptotic tests of ?, ? and ?. We also compare with the model-X knockoff procedures of ? and ? for variable selection with false discovery control. Our test is found to be as statistically powerful as the alternatives, with minimal false discovery, while reducing computational costs by several orders of magnitude. Moreover, we show that our feature importance measure is robust to noise in the data, delivering stable feature importance rankings. This contrasts sharply with widely-used algorithmic feature importance measures including SHAP (?) which are found to be highly sensitive to noise in the data, leading to unstable feature rankings that make it difficult to draw reliable inferences regarding variable importance. Our findings complement the discussion of the theoretical caveats of algorithmic importance measures in ?, ? and others.
@@
Empirical analyses demonstrate the ability of our framework to inform high-stakes decisions with legal, economic, and regulatory implications. In the context of a credit scoring task using large ensembles of gradient boosting machines, our approach identifies and ranks several significant features describing past payment behavior and social factors. Interestingly, we find that gender, marital status, and age of card holders (all of which are protected characteristics under US lending laws) are significant predictors of card default, after controlling for past payment behavior. In the context of a mortgage state transition prediction task using a recently proposed set-sequence model based on long-convolution architectures (?) and trained on extensive panel data, our test identifies time-varying macro-economic variables such as mortgage rates, as well as features summarizing past state transitions as significant predictors of borrower behavior.
@@
The rest of this introduction discusses the related literature. Section 2 introduces the notion of feature effects while Section 3 develops the testing procedure. Section 4 analyzes the median feature effect as an importance measure dual to the test. Section 5 studies the performance of the test and importance measure in synthetic regression and classification settings. Section 6 deploys the test on two real-world data sets. Section 7 concludes. The proofs of technical results are given in an appendix. A Python package implementing our testing and importance procedures is available at https://github.com/ji-chartsiri/AICO.
@@ 1.1 Related literature
There is a large body of work developing algorithmic feature importance measures; see ? for an overview. Popular examples include measures based on Shapley values (e.g. SHAP, see ? and many others), measures based on local approximations (e.g. LIME, see ?), and measures obtained by permutation methods (e.g. ?, ?). These measures provide valuable insight into the role of features but unlike the feature importance measures delivered by our hypothesis testing approach, they typically do not offer statistical guarantees or $p$-values. We demonstrate that algorithmic measures can be highly susceptible to noise in the data so the lack of confidence intervals is concerning. Moreover, algorithmic measures such as SHAP can be biased in the presence of feature correlation (?). Our approach can deliver importance measures that are robust to feature correlation. Finally, we demonstrate that our method achieves orders-of-magnitude improvements in computational efficiency over widely used algorithmic importance approaches, such as SHAP and permutation scores.
@@
The statistical and econometric literature has developed significance tests for the variables in nonparametric regression models. ?, ?, ?, ?, ? and others employ kernel methods to construct asymptotic-based testing procedures. ?, ? and ? and others construct gradient-based asymptotic tests using neural networks. Unlike these tests, ours are model-agnostic, non-asymptotic, deliver exact $p$-values and confidence intervals, and can accommodate a wide range of data structures.
@@
Pursuing a model-agnostic approach, ?, ?, ?, ?, ?, ? and others develop non-asymptotic tests for a conditional independence hypothesis of feature significance in the model-X knockoff setting of ?. [^2] Conditional independence testing is challenging: ? show that no test of conditional independence can be uniformly valid against arbitrary alternatives. We show that, for an appropriate choice of the masking baseline, our null is generally weaker than the conditional independence null studied in these papers; equivalence holds under additional conditions. ? construct non-asymptotic tests based on the drop in model predictive accuracy due to removing features (LOCO: Leave-out-covariates). We, in contrast, test the improvement in model score due to masking a feature.
@@
The computational requirements of the aforementioned model-agnostic testing procedures are a limiting factor. The LOCO test requires re-training the model for each feature to be tested, which can be prohibitive for feature-rich models that are costly to train. The model-X knockoff procedures require constructing an auxiliary model of the joint feature distribution, which can be as challenging and computationally costly as constructing the model to be tested. In fact, ? show that no algorithm can efficiently compute nontrivial knockoffs for arbitrary feature distributions. Our approach does not require any model re-training or the construction of auxiliary models, delivering an unprecedented degree of computational efficiency. Our numerical results demonstrate that the drastic gains in computational performance we achieve do not come at the expense of statistical performance.
@@ 2 Feature effect
We consider data $Z_{1},\ldots,Z_{m}$ with $Z_{i}=(X_{i},Y_{i})$. Here, $X_{i}=(X^{1}_{i},\ldots,X^{d}_{i})$ is a $d$-dimensional feature vector taking values in a set $\mathcal{X}$ and $Y_{i}$ is a response taking values in a set $\mathcal{Y}$. In a typical regression setting with tabular data, $\mathcal{X}\subseteq\mathbb{R}^{d}$ and $\mathcal{Y}=\mathbb{R}$. We make no assumptions on the distribution of the $Z_{i}$’s.
@@
We take as given a random split of $m]=\{1,\ldots,m\}$ into non-intersecting subsets $I_{1}$ and $I_{2}$, which are used for training/validation and testing purposes, respectively. This partitioning is standard in supervised learning settings. In our testing framework, data splitting facilitates post-selection inference.
@@
We also take as given a model (algorithm) $f:\mathcal{X}\to\mathcal{M}$ trained on $\{Z_{i}\}_{i\in I_{1}}$, which is to be tested. We do not make any assumptions on the form of the modeling target; $f$ could represent $\mathbb{E}(Y_{i}|X_{i})$, a conditional quantile, or some other statistic of $Y_{i}$. We also do not impose assumptions on the training criterion (eg. $L_{2}$ error) or approach (eg. empirical risk minimization) being used for training.
@@
We would like to test the statistical significance of a feature $\ell\in[d]$ under the model $f$. (Unless noted otherwise, $\ell$ is fixed.) In the special case that $f$ is linear, standard asymptotic procedures such as the $t$-test are available. We develop a simple, non-asymptotic and model-agnostic testing procedure that does not require a special model structure, and is valid much more broadly for models $f$ of essentially any form.
@@
We consider the impact on model performance of a feature’s variability across the test set $I_{2}$. Performance is measured by a deterministic score function $S:\mathcal{M}\times\mathcal{Y}\to\mathbb{R}$, which relates model output to response. Specifically, for a given feature value $x\in\mathcal{X}$ and response value $y\in\mathcal{Y}$, the score $S(f(x),y)$ evaluates the quality of the model output $f(x)$ in view of $y$. Higher scores reflect better performance. A natural choice for $S$ is the negative of the loss function used for the training of $f$. Examples include the negative $L_{2}$ loss $S(m,y)=-(m-y)^{2}$ in a regression context, the negative pinball loss in a quantile regression context, or the negative (binary or multi-class) cross-entropy loss in a classification context. With these specifications $S$ measures predictive accuracy. Other specifications of $S$ can measure model performance (e.g. revenue) in the context of a downstream screening or decisioning application of $f$, for example the financial loss caused by a poor decision due to mis-classification. Finally, we can specify a function $S$ that does not depend on the second argument representing the response, to measure the behavior (e.g. sensitivity) of the model output on its own, without reference to the response.
@@
We evaluate the effect on model score of a feature’s variability across the test set. For each $i\in I_{2}$ we construct from $X_{i}$ a baseline feature value $\underline{X}_{i}\in\mathcal{X}$ serving as a reference and a comparison feature value $\overline{X}_{i}\in\mathcal{X}$. In the baseline, the value of the feature to be tested is masked with a reference value whose choice is further discussed below. The comparison value $\overline{X}_{i}$ is unmasked so that the predicted response $f(\overline{X}_{i})$ can harness the actual sample $i$ value of the feature to be tested. We propose to measure the feature effect by the resulting score improvement.
@@
For a given score function $S$, baseline feature value $\underline{X}_{i}$ and comparison feature value $\overline{X}_{i}$, the feature effect on test sample $i\in I_{2}$ is the score delta
@@
Note that masking feature values for purposes of measuring a feature effect has significant advantages over alternative approaches such as randomly shuffling feature values (e.g. ? and ?) or dropping a feature from the feature vector. Shuffling introduces excess variance into the measurement procedure and dropping features requires training sub-models, which (drastically) reduces the computational efficiency of the measurement procedure. Our masking approach avoids these issues.
@@ 2.1 Baseline and comparison features
Along with the selection of the score function $S$, the selection of baseline and comparison feature values enables users to tailor the feature effect to the model and data structures at hand as well as the testing objectives. We provide a conceptional framework and several examples. Throughout, fix a feature $\ell\in[d]$ that is to be tested; our notation usually suppresses the dependence on $\ell$. We start by selecting a non-empty set $\underline{I}\subseteq I_{1}$ of training samples that will be used to construct the baseline $\underline{X}_{i}$ for the test sample $i\in I_{2}$. In a second step, we aggregate the samples in $\underline{I}$ to obtain $\underline{X}_{i}$. Alternative aggregators include the mean, a quantile or a mode of $\{X_{k}:k\in\underline{I}\}$.
@@
Take $\underline{I}=I_{1}$ and let $\underline{X}_{i}=(v^{1},\ldots,v^{d})$ for all $i\in I_{2}$ where $v^{j}$ is the mean of $\{X^{j}_{k}:k\in\underline{I}\}$ for each $j\in[d]$. The predicted response $f(\underline{X}_{i})$ is constant across the test set. If feature $\ell=1$ is to be tested, select the comparison value $\overline{X}_{i}=(X_{i}^{1},v^{2},\ldots,v^{d})$. The predicted response $f(\overline{X}_{i})$ harnesses feature 1’s actual sample value, $X_{i}^{1}$. The feature effect $\Delta_{i}$ represents the associated score improvement.
An alternative choice is $\underline{X}_{i}=(v^{1},X^{2}_{i},\ldots,X^{d}_{i})$ and $\overline{X}_{i}=X_{i}=(X_{i}^{1},\ldots,X^{d}_{i})$, again supposing feature $\ell=1$ is to be tested. The predicted baseline response $f(\underline{X}_{i})$ is no longer constant across the test set, and accounts for the variation of features $X^{2}_{i},\ldots,X^{d}_{i}$. Here $\Delta_{i}$ measures a conditional feature effect.
Choices other than the mean may be appropriate for $v^{j}$. For example, in the case of a one-hot feature $X^{j}_{i}\in\{0,1\}$ which might encode a categorical variable, sensible choices for $v^{j}$ are a mode of $\{X^{j}_{k}:k\in\underline{I}\}$ or the value $1-X^{j}_{i}$. [^3] See ? for extensions.
Consider choices of $\underline{I}$ beyond the entire training set $I_{1}$. We can pick specific samples from $I_{1}$ by first fixing a reference point $r$ in $\mathcal{Y}$, $\mathcal{M}$, or $\mathcal{X}$ and then selecting the samples from $I_{1}$ that lie in or outside a specified neighborhood of $r$. The point $r$ can represent the (weighted) mean, mode or a $\lambda$-quantile of the responses $\{Y_{k}:k\in I_{1}\}$, the predicted responses $\{f(X_{k}):k\in I_{1}\}$, or the values of the feature $\ell$ to be tested, $\{X^{\ell}_{k}:k\in I_{1}\}$. Potentially interesting special cases are quantiles for $\lambda\in\{0,1\}$. Concretely, let $Y_{r}\in\mathcal{Y}$ be a reference response value and $d$ be a distance measure on $\mathcal{Y}$. We can now take $\underline{I}$ as the set of samples $k\in I_{1}$ for which $d(Y_{k},Y_{r})$ does not exceed or surpasses some threshold value. Other criteria can be envisioned, for example samples $k\in I_{1}$ for which $Y_{k}$ exceeds or falls short of $Y_{r}$. A similar approach applies to reference points in $\mathcal{M}$ or $\mathcal{X}$.
@@
Consider the reference point $\min_{k\in I_{1}}f(Y_{k})$ in $\mathcal{M}$, which is a quantile of the predicted responses $\{f(Y_{k}):k\in I_{1}\}$ at level $\lambda=0$. For example, in a binary soft classification setting such as fraud detection or credit scoring where $Y_{i}=0$ represents the preferred (beneficial) outcome, this reference point represents the training samples with minimum predicted risk. [^4] See ?. We can take $\underline{I}=\{\mbox{argmin}_{k\in I_{1}}f(Y_{k})\}$ as the set of these samples and define $\underline{X}_{i}$ and $\overline{X}_{i}$ as in Example 2.1. The associated feature effect $\Delta_{i}$ represents the score delta relative to the minimum risk baseline. It measures how much excess risk a feature explains.
A few basic principles guide the selection of reference points and the construction of the set $\underline{I}\subseteq I_{1}$. We wish to obtain values $\underline{X}_{i}$ and $\overline{X}_{i}$ that are sufficiently distinct, to avoid feature effects $\Delta_{i}$ that are biased low which result in significance tests with low power.555? discuss these issues in the context of image classification. Similarly, we want to avoid saturation, where the predicted response is relatively flat in the neighborhood of $\underline{X}_{i}$. Finally, we might want to avoid the selection of values $(\underline{X}_{i},\overline{X}_{i})$ that are off the data manifold or in low-density areas of the training space. It is important to select an appropriate reference point and rule used to aggregate across $\underline{I}$ (see Example 2.1). Alternatively, we might select an appropriate baseline value from the set $\{X_{k}:k\in I_{1}\}$, namely an element that is near the raw off-manifold value emerging from the baseline selection procedure. [^6] We also allow for the selection and aggregation of alternative baselines for a given sample. A general tool is sample-dependent or feature-dependent selection of $(\underline{X}_{i},\overline{X}_{i})$.

### v4

@@ Abstract
Machine learning has become a central tool across scientific, industrial, and policy domains. Algorithms now identify chemical properties, forecast disease risk, screen borrowers, and guide public interventions. Yet this predictive power often comes at the cost of transparency: we rarely know which input features truly drive a model’s predictions. Without such understanding, researchers cannot draw reliable scientific conclusions, practitioners cannot ensure fairness or accountability, and policy makers cannot trust or govern model-based decisions. Despite its importance, existing tools for assessing feature influence are limited—most lack statistical guarantees, and many require costly retraining or surrogate modeling, making them impractical for large modern models. We introduce AICO, a broadly applicable framework that turns model interpretability into an efficient statistical exercise. AICO asks, for any trained regression or classification model, whether each feature genuinely improves model performance. It does so by masking the feature’s information and measuring the resulting change in performance. The method delivers exact, finite-sample inference—exact feature $p$-values and confidence intervals—without any retraining, surrogate modeling, or distributional assumptions, making it feasible for today’s large-scale algorithms. In both controlled experiments and real applications—from credit scoring to mortgage-behavior prediction—AICO consistently pinpoints the variables that drive model behavior, providing a fast and reliable path toward transparent and trustworthy machine learning.
@@ 1 Introduction
Machine learning has become integral to modern science, industry, and public policy. Algorithms now predict molecular activity, identify disease markers, forecast financial defaults, and guide the allocation of social resources. Yet even as these systems achieve extraordinary predictive accuracy, their inner workings remain largely opaque. We often do not know which input features truly shape a model’s predictions, which merely correlate by chance, or which may encode unwanted biases. This opacity limits scientific understanding, hinders regulatory oversight, and undermines public trust in data-driven decision making. Despite extensive progress in explainable artificial intelligence, most tools for measuring feature importance lack statistical guarantees or depend on computationally costly retraining or surrogate modeling procedures, restricting their use in today’s large, complex models. As a result, reconciling interpretability with scalable inference remains an open challenge for modern machine-learning practice.
We introduce AICO, a tractable method for feature significance testing in any supervised learning model, whether for regression or classification. AICO transforms the question of interpretability into one of formal statistical inference: it tests whether a feature contributes meaningfully to a model’s predictive or decision-making performance, using only the model’s fitted predictions and observed outcomes. AICO constructs test statistics based on the change in model performance when a feature is masked. It is model- and distribution-agnostic: it makes no assumptions about the model class, data distribution, or estimand (for example, conditional mean or quantile). The method does not rely on retraining or surrogate models, making it well suited for analyzing complex, high-dimensional algorithms that are costly to train. It bridges a long-standing gap between machine-learning practice and classical statistical inference, offering a tractable route to transparent, trustworthy prediction in today’s large-scale applications.
@@
The AICO feature effect—the change in model performance due to masking—quantifies a feature’s importance for each individual sample. For a null feature, these effects form a distribution across test samples whose median is non-positive. A uniformly most powerful randomized sign test achieves finite-sample exactness: its size (type-I error probability) exactly matches the nominal level (e.g., $5\%$). The procedure produces exact, randomized feature $p$-values whose conditional distribution, given the data, is uniform on an interval that can be reported in practice. It is simple, robust, and computationally efficient, relying on minimal assumptions and remaining valid under broad conditions. These properties enable valid inference on a model’s features without relying on large-sample approximations or parametric assumptions.
@@
The median feature effect being tested provides a natural, population-level measure of feature importance. We develop randomized, exact-coverage confidence intervals for this median that, together with the $p$-values, enable principled inference on a feature’s influence. The availability of confidence intervals distinguishes the median from widely used algorithmic measures of feature importance (see Section 1.1). The median also has several desirable properties that depend on the choice of masking value. It can be robust to correlation among features—a pervasive challenge in practical data settings—and, in linear regression, it is proportional to the square of the feature’s regression coefficient, the canonical measure of importance in the linear context. Finally, we provide conditions under which a feature’s median equal to zero is equivalent to conditional independence of that feature and the response given the remaining covariates.
@@
Extensive experiments on synthetic “known-truth” tasks demonstrate the statistical and computational performance of our approach in nonlinear regression and classification settings. We compare AICO with powerful, model-agnostic, and non-asymptotic tests by ?, ?, and ?, as well as with the model-X knockoff procedures of ? and ? for variable selection under false discovery control. Our test achieves comparable or higher statistical power than these alternatives, with minimal false discoveries, while reducing computational cost by several orders of magnitude. We further show that the AICO feature-importance measure is robust to noise in the data, yielding stable feature rankings. This stands in sharp contrast to widely used algorithmic importance measures such as SHAP (?), which are highly sensitive to data perturbations and can produce unstable rankings that hinder reliable inference about variable importance. These empirical findings reinforce recent theoretical discussions of the limitations of algorithmic importance measures in ?, ?, and related work.
@@
Empirical analyses demonstrate the ability of our framework to inform high-stakes decisions with legal, economic, and regulatory implications. In a credit scoring task using ensembles of gradient-boosting models, AICO identifies and ranks several statistically significant features describing past payment behavior and social factors. Notably, the analysis reveals that gender, marital status, and age of card holders—protected characteristics under U.S. lending laws—are statistically significant predictors of card default, even after controlling for past payment behavior. In a mortgage state–transition prediction task based on a recently proposed set–sequence model with long-convolution architectures (?) and trained on extensive panel data, AICO detects time-varying macroeconomic variables such as mortgage rates, as well as features summarizing past state transitions, as significant predictors of borrower behavior.
@@
The remainder of this introduction discusses the related literature. Section 2 introduces the notion of feature effects, and Section 3 develops the testing procedure. Section 4 analyzes the median feature effect as an importance measure dual to the test. Section 5 examines the performance of the test and importance measure in synthetic regression and classification settings, while Section 6 applies the method to two real-world data sets. Section 7 concludes. Proofs of the theoretical results are provided in an Appendix. An open-source Python package implementing the testing and importance procedures is available at https://github.com/ji-chartsiri/AICO.
@@ 1.1 Related literature
There is a large body of work developing algorithmic feature importance measures; see ? for an overview and ? for a unifying framework for interpretability methods. Popular examples include measures based on Shapley values (e.g. SHAP, see ? and many others), measures based on local approximations (e.g. LIME, see ?), and measures obtained by permutation methods (e.g. ?, ?). These measures provide valuable insight into the role of features but unlike the feature importance measures delivered by our hypothesis testing approach, they typically do not offer statistical guarantees or $p$-values. We demonstrate that algorithmic measures can be highly susceptible to noise in the data so the lack of confidence intervals is concerning. Moreover, algorithmic measures are often biased in the presence of feature correlation (?, ?). Our approach can deliver importance measures that are robust to feature correlation. Finally, we demonstrate that our method achieves orders-of-magnitude improvements in computational efficiency over widely used algorithmic importance approaches, such as SHAP and permutation scores.
@@
The statistical and econometric literature has long studied significance testing for variables in nonparametric regression models. ?, ?, ?, ?, ?, and others employ kernel-based methods to construct asymptotic testing procedures. ?, ?, ?, and others develop gradient-based asymptotic tests using neural networks. ? develop permutation-based asymptotic tests for random forests; see also ? for an alternative approach using $U$-statistics. In contrast, the AICO tests are model-agnostic and non-asymptotic, deliver exact $p$-values and confidence intervals, and accommodate a broad range of data structures.
@@
Pursuing a model-agnostic approach, ?, ?, ?, ?, ?, ?, and others develop non-asymptotic tests for conditional independence in the model-X knockoff framework of ?. Conditional independence testing is inherently difficult: ? show that no such test can be uniformly valid against arbitrary alternatives. We show that, for an appropriate choice of the feature masking value, the AICO null is generally weaker than the conditional-independence null studied in these papers, with equivalence holding under additional conditions. ? suggest model-agnostic tests based on the drop in model predictive accuracy when features are removed from the model (LOCO: leave-out-covariates; see also ? and ?). In contrast, AICO tests the change in model performance due to masking a feature.
@@
The computational demands of the aforementioned model-agnostic testing procedures are a major limiting factor. The LOCO test requires retraining the model for each feature to be tested, which can be prohibitive for feature-rich models that are costly to train. The model-X knockoff procedures require constructing an auxiliary model of the joint feature distribution, a task that can be as challenging and computationally expensive as fitting the predictive model itself. Indeed, ? show that no algorithm can efficiently compute nontrivial knockoffs for arbitrary feature distributions. In contrast, AICO requires neither model retraining nor auxiliary modeling, delivering an unprecedented degree of computational efficiency. Our numerical experiments demonstrate that these drastic gains in computational speed do not come at the expense of statistical power or inferential accuracy.
@@ 2 Feature effect
Consider data $Z_{1},\ldots,Z_{m}$ with $Z_{i}=(X_{i},Y_{i})$. Here, $X_{i}=(X^{1}_{i},\ldots,X^{d}_{i})$ is a feature vector taking values in a set $\mathcal{X}$, and $Y_{i}$ is a response taking values in a set $\mathcal{Y}$. For example, in a typical regression setting with tabular data, $\mathcal{X}\subseteq\mathbb{R}^{d}$ and $\mathcal{Y}=\mathbb{R}$. We make no assumptions on the distribution of the $Z_{i}$’s.
@@
We take as given a random split of $m]=\{1,\ldots,m\}$ into non-intersecting subsets $I_{1}$ and $I_{2}$, used for training and testing, respectively. Such data splitting is standard in supervised learning and, in our setting, facilitates post-selection inference.
@@
We also take as given a model (algorithm) $f:\mathcal{X}\to\mathcal{M}$ trained on $\{Z_{i}\}_{i\in I_{1}}$, which is to be tested. We make no assumptions on the form of the estimand; $f$ may represent the conditional mean, a quantile, or another statistic of $Y_{i}$. We likewise do not impose assumptions on the training criterion (e.g., $L_{2}$ error) or on the optimization approach (e.g., empirical risk minimization) used to fit $f$.
@@
We aim to test the statistical significance of a feature $\ell\in[d]$ under the fitted model $f$. (Unless noted otherwise, $\ell$ is fixed.) When $f$ is linear, classical asymptotic procedures such as the $t$-test are available. We develop instead a simple, non-asymptotic, and model-agnostic testing procedure that does not rely on a specific model structure and remains valid for models $f$ of essentially any form.
@@
We consider the impact of a feature’s variability on model performance across the test set $I_{2}$. Performance is measured by a deterministic score function $S:\mathcal{M}\times\mathcal{Y}\to\mathbb{R}$ that relates model output to response. For a given feature value $x\in\mathcal{X}$ and response value $y\in\mathcal{Y}$, the score $S(f(x),y)$ evaluates the quality of the model output $f(x)$ relative to $y$, with higher scores indicating better performance. A natural choice for $S$ is the negative of the loss function used to train $f$. Examples include the negative $L_{2}$ loss $S(m,y)=-(m-y)^{2}$ in regression, the negative pinball loss in quantile regression, or the negative (binary or multi-class) cross-entropy loss in classification. With these specifications, $S$ measures predictive accuracy. Other choices of $S$ can quantify model performance in downstream screening or decisioning applications of $f$ (e.g., revenue or cost), such as the financial loss caused by a poor decision due to misclassification. Finally, one may specify a function $S$ that does not depend on the second argument $y$ to measure the behavior of the model output itself (for example, its sensitivity to the inputs).
@@
We evaluate the effect of a feature’s variability on the model score across the test set. For each $i\in I_{2}$, we construct from $X_{i}$ a baseline feature value $\underline{X}_{i}\in\mathcal{X}$, serving as a reference, and a comparison feature value $\overline{X}_{i}\in\mathcal{X}$. In the baseline, the value of the feature to be tested is masked with a reference value, the choice of which is discussed further below. The comparison value $\overline{X}_{i}$ is unmasked so that the predicted response $f(\overline{X}_{i})$ can leverage the actual sample-$i$ value of the feature being tested. We measure the feature’s effect by the corresponding improvement in model score.
@@
For a given score function $S$, baseline feature value $\underline{X}_{i}$, and comparison feature value $\overline{X}_{i}$, the feature effect on test sample $i\in I_{2}$ is defined as the score difference
@@
Masking feature values to measure a feature’s effect offers significant advantages over alternatives such as random shuffling, random perturbation or dropping the feature from the vector. Random procedures introduce additional variance into the measurement, while dropping features requires training sub-models, which drastically reduces computational efficiency. Our masking approach avoids both issues, yielding stable and computationally efficient estimates of feature importance.
@@ 2.1 Baseline and comparison features
Along with the selection of the score function $S$, the choice of baseline and comparison feature values allows users to tailor the feature effect to the model, data structure, and testing objectives at hand. We provide a conceptual framework and several examples. Throughout, fix a feature $\ell\in[d]$ to be tested; our notation usually suppresses the dependence on $\ell$.
We begin by selecting a non-empty set $\underline{I}\subseteq I_{1}$ of training samples that will be used to construct the baseline $\underline{X}_{i}$ for each test sample $i\in I_{2}$. In a second step, we aggregate the samples in $\underline{I}$ to obtain $\underline{X}_{i}$. Possible aggregation rules include the mean, a quantile, or a mode of $\{X_{k}:k\in\underline{I}\}$.
@@
Take $\underline{I}=I_{1}$ and let $\underline{X}_{i}=(v^{1},\ldots,v^{d})$ for all $i\in I_{2}$, where $v^{j}$ is the mean of $\{X^{j}_{k}:k\in\underline{I}\}$ for each $j\in[d]$. Then $f(\underline{X}_{i})$ is constant across the test set. For $\ell=1$, take $\overline{X}_{i}=(X_{i}^{1},v^{2},\ldots,v^{d})$. Here, the feature’s actual sample value $X_{i}^{1}$ is added back in while all other components remain fixed—an operation that gives rise to the name AICO (Add-In COvariates). The predicted response $f(\overline{X}_{i})$ therefore leverages the true variation in feature 1, and the feature effect $\Delta_{i}$ represents the resulting improvement in model score.
An alternative choice is $\underline{X}_{i}=(v^{1},X_{i}^{2},\ldots,X_{i}^{d})$ and $\overline{X}_{i}=X_{i}=(X_{i}^{1},\ldots,X_{i}^{d})$, again with $\ell=1$. Here, the predicted baseline response $f(\underline{X}_{i})$ is no longer constant across the test set, as it incorporates variation in the remaining features $X_{i}^{2},\ldots,X_{i}^{d}$. Adding the true value $X_{i}^{1}$ back into this context performs the same AICO operation, now yielding a conditional feature effect $\Delta_{i}$ that measures the incremental contribution of feature 1 given the others.
Choices other than the mean may also be appropriate for $v^{j}$. For instance, when $X_{i}^{j}\in\{0,1\}$ encodes a categorical variable, the mean of $\{X^{j}_{k}:k\in\underline{I}\}$ may not represent a valid feature value. Sensible alternatives include the mode of $\{X^{j}_{k}:k\in\underline{I}\}$ or the complementary value $1-X_{i}^{j}$.
The set $\underline{I}$ need not equal the entire training set $I_{1}$. We can select specific samples from $I_{1}$ by first fixing a reference point $r$ in $\mathcal{Y}$, $\mathcal{M}$, or $\mathcal{X}$ and then choosing the samples from $I_{1}$ that lie in or outside a neighborhood of $r$. The point $r$ can represent, for example, the (weighted) mean, mode, or a $\lambda$-quantile of the responses $\{Y_{k}:k\in I_{1}\}$, of the predicted responses $\{f(X_{k}):k\in I_{1}\}$, or of the feature values $\{X^{\ell}_{k}:k\in I_{1}\}$ for the feature being tested. Quantiles with $\lambda\in\{0,1\}$ are particularly interesting special cases. Concretely, let $Y_{r}\in\mathcal{Y}$ be a reference response value and $d$ a distance measure on $\mathcal{Y}$. We can then take $\underline{I}$ as the set of samples $k\in I_{1}$ for which $d(Y_{k},Y_{r})$ does not exceed (or surpasses) a given threshold. Other criteria can also be used, such as selecting samples for which $Y_{k}$ exceeds or falls short of $Y_{r}$. A similar approach applies when $r$ is a reference point in $\mathcal{M}$ or $\mathcal{X}$.
@@
Consider the reference point $\min_{k\in I_{1}}f(Y_{k})$ in $\mathcal{M}$, which is the $\lambda=0$ quantile of the predicted responses $\{f(Y_{k}):k\in I_{1}\}$. In a binary soft-classification setting such as fraud detection or credit scoring, where $Y_{i}=0$ denotes the preferred (beneficial) outcome, this reference point corresponds to the training samples with minimum predicted risk. [^2] See ?. We can take $\underline{I}=\{\mbox{argmin}_{k\in I_{1}}f(Y_{k})\}$ as the set of these samples and define $\underline{X}_{i}$ and $\overline{X}_{i}$ as in Example 2.1. The resulting feature effect $\Delta_{i}$ represents the score delta relative to the minimum-risk baseline and measures how much excess risk a feature explains.
A few general principles guide the selection of reference points and the construction of $\underline{I}\subseteq I_{1}$. We aim to obtain values $\underline{X}_{i}$ and $\overline{X}_{i}$ that are sufficiently distinct, avoiding feature effects $\Delta_{i}$ that are biased toward zero and lead to significance tests with low power.333? discuss related issues in the context of image classification. At the same time, we seek to avoid saturation, where the predicted response is relatively flat in the neighborhood of $\underline{X}_{i}$. Further, we might want to avoid the selection of values $(\underline{X}_{i},\overline{X}_{i})$ that are off the data manifold or in low-density areas of the training space. It is important to choose an appropriate reference point and aggregation rule for $\underline{I}$ (see Example 2.1). Alternatively, one may select a suitable baseline value directly from the set $\{X_{k}:k\in I_{1}\}$—namely, an element near the raw off-manifold value produced by the baseline selection procedure. [^4] We also allow for the selection and aggregation of alternative baselines for a given sample. A general approach is sample-dependent or feature-dependent selection of $(\underline{X}_{i},\overline{X}_{i})$.